OpenStudy (anonymous):

The price of 9-volt batteries is increasing according to the function below, where t is years after January 1, 1980. During what year will the price reach $5? P(t) = 1.1 * e^0.047t

OpenStudy (anonymous):

5=1.1* e^0.047 t Solve for t

OpenStudy (anonymous):

5/1.1=e^0.047 t t = ln(5/11)/0.047